# WeChat-YATT: A Simple, Scalable and Balanced RLHF Trainer

### v1

@@ 6 Conclusion
In this work, we presented WeChat-YATT, a simple, scalable, and balanced RLHF training framework designed to address key bottlenecks in large-scale, multi-model reinforcement learning workflows. By introducing a parallel controller programming model and a dynamic scaling placement schema, WeChat-YATT effectively mitigates memory and communication bottlenecks, and adapts to dynamic workloads with improved hardware utilization. Our design emphasizes practicality and flexibility, aiming to reduce engineering complexity while supporting efficient orchestration of complex RLHF pipelines.
WeChat-YATT has already produced model that deployed in WeChat’s production environment, supporting real-world business models that serve the entire user base. This practical deployment demonstrates the framework’s robustness and effectiveness at scale. While there remain open challenges in further optimizing resource allocation and supporting increasingly diverse model architectures, we believe WeChat-YATT provides a solid foundation for future research and practical deployment of large-scale, human-aligned models.

### v3

@@ 6 Conclusion
In this work, we present WeChat-YATT, a simple, scalable, and balanced RLHF training framework that addresses critical bottlenecks in large-scale, multi-model reinforcement learning workflows. By introducing a parallel controller programming model and a dynamic scaling placement schema, WeChat-YATT effectively mitigates memory and communication bottlenecks, and adapts to dynamic workloads with improved hardware utilization. Our design emphasizes practicality and flexibility, aiming to reduce engineering complexity while supporting efficient orchestration of complex RLHF pipelines.
WeChat-YATT has already produced various models that deployed in WeChat’s production environment, supporting real-world business applications that serve the entire user base. This practical deployment demonstrates the framework’s robustness and scalability. Although challenges remain in optimizing resource allocation and accommodating increasingly diverse model architectures, we believe WeChat-YATT offers a strong foundation for advancing research and deploying large-scale, human-aligned models in practice.